# Quantifying Observational Projection Effects with a Simulation-based hot CGM model

### v1

@@ 5 Discussion
This work uses a TNG300-based forward model for the hot CGM emission and presents, for the first time, the effect of locally correlated large-scale structure around a halo and the effect of misclassified centrals in stacked hot CGM galactocentric profiles. Our findings are vital in light of our explorations for studying X-ray emission around lower mass galaxies with stacking experiments. We divide the discussion of our results by focusing on the projection effects due to the locally correlated environment, first, in Sec. 5.1 and second, the effect of misclassified centrals in Sec. 5.2.
@@ 5.2 The Effect of Misclassified Centrals
We quantify the impact of the misclassified centrals (satellite-boost) residing in the same stellar mass bin as the central galaxies in Fig. 78, and 9 for the stellar mass bins $10^{10.5-11}$, $10^{11-11.2}$, and $10^{11.2-11.5}$, respectively. For satellite galaxies with stellar masses of MW-like galaxies, the emission around satellites dominates over central emission because the satellites of the stellar mass range $10^{10.5-11}{\rm M}_{\odot}$ reside in distinct halos with masses $M_{\rm 200m}=10^{11.9-15.1}$, corresponding to a mean halo mass $M_{200m}=10^{13.7}\ {\rm M}_{\odot}$. We find that the satellite subhalos probe the emission of their parent halos, contributing to the stack with off-centred surface brightness profiles of their more massive parent halo. This contribution causes shallower slopes for the total emission (central and satellites) in the MW-stellar mass bin. In cases where $30\%$, $10\%$, or $1\%$ of the satellites contribute to the total emission, we have shown that they dominate the measured total profile at radii $\geq 0.06\times R_{200m}$, $\geq 0.2\times R_{200m}$, and $\geq 0.3\times R_{200m}$. We further investigate the effect of misclassified centrals in the more massive stellar mass bins of $10^{11-11.2}\ {\rm M}_{\odot}$, and $10^{11.2-11.5}\ {\rm M}_{\odot}$ in Fig. 8 and Fig. 9, respectively. We find the effect of misclassified centrals, with contamination fractions of $0.01$, $0.1$, and $0.3$, less affects the $10^{11-11.2}\ {\rm M}_{\odot}$, and $10^{11.2-11.5}\ {\rm M}_{\odot}$ mass bins as opposed to the MW-stellar mass bin. We reaffirm this with the trend shown in Fig. 10, where the fraction of intrinsic emission at the mean $R_{500c}$ is lowest for the MW stellar mass bin for all contamination levels, as opposed to the more massive stellar mass bins.
@@
To understand why the MW-stellar mass bin is most affected by the effect of misclassified centrals, we first look at the mean halo masses of the satellites in the higher stellar mass bins of $10^{11-11.2}\ {\rm M}_{\odot}$ and $10^{11.2-11.5}\ {\rm M}_{\odot}$. The satellites reside in mean distinct halos of $\sim 10^{13.8}\ {\rm M}_{\odot}$ and $\sim 10^{13.9}\ {\rm M}_{\odot}$, respectively. This is a direct implication arising from the difference in the stellar-to-halo-mass relation for central and satellite galaxies (Shuntov et al. 2022). Given that the mean halo mass of the central galaxies in the LC-TNG300 stellar mass bins of $10^{11-11.2}\ {\rm M}_{\odot}$, and $10^{11.2-11.5}\ {\rm M}_{\odot}$ is $10^{13.3}\ {\rm M}_{\odot}$ and $10^{13.6}\ {\rm M}_{\odot}$, respectively, we require large fraction of contamination to see the effect of misclassified centrals compared to that for the MW-stellar mass bin. More precisely, we find that the effect of misclassified centrals dominates over the central galaxy emission at $R_{500c}$ with satellite contamination fractions, $f_{\rm cont,\ sat}$, of $0.02,0.28$, and $0.72$ for the stellar mass bins $10^{10.5-11}\ {\rm M}_{\odot}$, $10^{11-11.25}\ {\rm M}_{\odot}$, and $10^{11.25-11.5}\ {\rm M}_{\odot}$, respectively. Hence, the contamination fractions probed here ($f_{\rm cont,\ sat}=0.01,0.1,$ and $0.3$), which are closer to the realistic $f_{\rm cont,\ sat}$ in observations, do not significantly affect the stellar mass bins $10^{11-11.25}$, and $10^{11.25-11.5}$, respectively.
@@
Given these findings, we conclude that the contribution of satellites in MW-like galaxy samples containing satellites must not be neglected in stacking analysis. Observationally, to disentangle the contamination of the misclassified centrals, we must fit the profiles contributing to the intrinsic central galaxy jointly with the component capturing the effect of misclassified centrals for a given fraction of satellite contamination. The analytical model encapsulating the central profile and the effect of misclassified centrals will be addressed in future work.
@@ 6 Summary
    We present an analytical model (Eq. 7) that well-describes the intrinsic $S_{X}$-profile in LC-TNG300 across the stellar mass bins $10^{10.5-11}{\rm M}_{\odot}$, $10^{11-11.2}{\rm M}_{\odot}$, and $10^{11.2-11.5}{\rm M}_{\odot}$ and halo mass bins of $10^{12.5-13}{\rm M}_{\odot}$, $10^{13-13.5}{\rm M}_{\odot}$, and $10^{13.5-14}{\rm M}_{\odot}$. We provide the best-fitting parameters for the analytical model in Tab. 4 and Tab. 4 for the stellar and halo mass bins explored in this work.
@@
    We present, for the first time, the effect of misclassified centrals in stacking experiments for three stellar mass bins: $10^{10.5-11}{\rm M}_{\odot}$, $10^{11-11.2}{\rm M}_{\odot}$, and $10^{11.2-11.5}{\rm M}_{\odot}$. We find that the contaminating effect of the misclassified centrals on the stacked profiles increases when the stellar mass decreases.
@@
    For the MW-like galaxies, we conclude that the contribution of satellites (or misclassified centrals) can not be neglected in stacking analysis (see Fig. 10). In cases where $30\%$, $10\%$, or $1\%$ of the satellites contribute to the total emission of MW-like galaxies, we have shown that they dominate the measured total $S_{X}$ profile at radii $\geq 0.06\times R_{200m}$, $\geq 0.2\times R_{200m}$, and $\geq 0.3\times R_{200m}$, respectively.
Modelling observed CGM profiles and comparing them with simulations is crucial to constrain the different galaxy formation models and to understand how the feedback and physics prescriptions affect the hot CGM profile. Current state-of-the-art cosmological hydrodynamical simulations are calibrated on the stellar mass function and successfully reproduce realistic galaxy populations. Despite this, Davies et al. (2020) show that EAGLE and IllustrisTNG predict different total gas mass fractions, affecting the observed CGM properties at MW-masses. Similarly, Khrykin et al. (2024) shows how the hot gas is sensitive to the different feedback variants within SIMBA. The model and methodology presented here provide the machinery to compare the measured hot CGM profiles among current simulations in future works.
Future X-ray missions, on the observation side, like Athena (Nandra et al. 2013), HUBS (Cui et al. 2020), LEM (Kraft et al. 2022) will push our current detection limits and provide us with (i) the spatial resolution to reach higher redshifts, (ii) better quantify point-like source contamination within the hot CGM, (iii) the spectral resolution to allow disentangling components via spectral fitting and (iv) the grasp to reach even fainter surface brightness levels. This work, focusing on modelling the projection effects, is a step towards exploiting the information provided by the next-generation telescopes to better understand the intrinsic hot CGM emission. The prospects of this framework would be to extend it to interpret hot X-ray CGM measurements in stacking experiments by accounting for all projection effects, i.e., not only the impact due to the local environment of the halo and the effect due to misclassified centrals but also the emission from other X-ray sources such as AGN and XRB.
@@ Acknowledgements.
SS would like to thank Fulvio Ferlito, Matteo Guardiani, Christian Garrel, Emre Bahar, and Jeremy Sanders for all the useful scientific discussions.
G.P. acknowledges financial support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program ”Hot Milk” (grant agreement No. 865637) and support from Bando per il Finanziamento della Ricerca Fondamentale 2022 dell’Istituto Nazionale di Astrofisica (INAF): GO Large program and from the Framework per l’Attrazione e il Rafforzamento delle Eccellenze (FARE) per la ricerca in Italia (R20L5S39T9).
Computations were performed on the HPC system Raven at the Max Planck Computing and Data Facility. We acknowledge the project support by the Max Planck Computing and Data Facility.

### v2

@@ 6 Discussion
This work uses a TNG300-based forward model for the hot CGM emission and presents, for the first time, the effect of locally correlated large-scale structure around a halo, the effect of misclassified centrals in stacked hot CGM galactocentric profiles and the effect of the choice of centre. Our findings are vital in light of our explorations for studying X-ray emission around lower mass galaxies with stacking experiments. We divide the discussion of our results by focusing on the projection effects due to the locally correlated environment, first, in Sec. 6.1, second, the effect of misclassified centrals in Sec. 6.2, and lastly, the effect of miscentering in Sec. 6.3.
@@ 6.1 Locally correlated environment
Figure 10: Fraction of X-ray emission from the intrinsic hot CGM for different levels of satellite contamination (misclassified centrals) in the total galaxy sample. The fraction of X-ray emission from the intrinsic hot CGM for misclassified central contamination fractions of $0.01$ (dotted line), $0.1$ (dash-dotted line), $0.3$ (dashed line), and $0.5$ (solid line) are shown for three different mass bins: MW-like galaxies with stellar-masses in $10^{10.5-11}\ {\rm M}_{\odot}$ (left panel), $10^{11-11.25}\ {\rm M}_{\odot}$ (middle panel), and $10^{11.25-11.5}\ {\rm M}_{\odot}$ (right panel). The shaded regions around each curve correspond to the uncertainty of the mean profile obtained by bootstrapping. The vertical black solid line at $242$ kpc in the left panel, $369$ kpc in the middle panel, and $484.5$ kpc in the right panel corresponds to the mean $R_{500c}$ of the respective mass bins with the shaded area signifying the minimum and maximum values. We find that the contaminating effect of the misclassified centrals on the stacked profiles increases when the stellar mass decreases.
@@ 6.2 The effect of misclassified centrals
We quantify the impact of the misclassified centrals (satellite-boost) residing in the same stellar mass bin as the central galaxies in Fig. 78, and 9 for the stellar mass bins $10^{10.5-11}$, $10^{11-11.25}$, and $10^{11.25-11.5}$, respectively. For satellite galaxies with stellar masses of MW-like galaxies, the emission around satellites dominates over central emission because the satellites of the stellar mass range $10^{10.5-11}{\rm M}_{\odot}$ reside in distinct halos with masses $M_{\rm 200m}=10^{12.3-15.1}$, corresponding to a mean halo mass $M_{200m}=10^{13.7}\ {\rm M}_{\odot}$. We find that the satellite subhalos probe the emission of their parent halos, contributing to the stack with off-centred surface brightness profiles of their more massive parent halo. This contribution causes shallower slopes for the total emission (central and satellites) in the MW-stellar mass bin. In cases where $30\%$, $10\%$, or $1\%$ of the satellites contribute to the total emission, we have shown that they dominate the measured total profile at radii $\geq 0.11\times R_{500c}$, $\geq 0.24\times R_{500c},$ and $\geq 1.04\times R_{500c}$, respectively. We further investigate the effect of misclassified centrals in the more massive stellar mass bins of $10^{11-11.25}\ {\rm M}_{\odot}$, and $10^{11.25-11.5}\ {\rm M}_{\odot}$ in Fig. 8 and Fig. 9, respectively. We find the effect of misclassified centrals, with contamination fractions of $0.01$, $0.1$, and $0.3$, less affects the $10^{11-11.25}\ {\rm M}_{\odot}$, and $10^{11.25-11.5}\ {\rm M}_{\odot}$ mass bins as opposed to the MW-stellar mass bin. We reaffirm this with the trend shown in Fig. 10, where the fraction of intrinsic emission at the mean $R_{500c}$ is lowest for the MW stellar mass bin for all contamination levels, as opposed to the more massive stellar mass bins.
@@
To understand why the MW-stellar mass bin is most affected by the effect of misclassified centrals, we first look at the host halo masses of the satellites in the higher stellar mass bins of $10^{11-11.25}\ {\rm M}_{\odot}$ and $10^{11.25-11.5}\ {\rm M}_{\odot}$. These satellites reside in distinct halos of mean $M_{\rm 200m}$ $\sim 10^{14}\ {\rm M}_{\odot}$ and $\sim 10^{14.2}\ {\rm M}_{\odot}$, respectively. This is a direct implication arising from the difference in the stellar-to-halo-mass relation for central and satellite galaxies (Shuntov et al. 2022). Given that the mean halo mass of the central galaxies in the LC-TNG300 stellar mass bins of $10^{11-11.25}\ {\rm M}_{\odot}$, and $10^{11.25-11.5}\ {\rm M}_{\odot}$ is $10^{13.3}\ {\rm M}_{\odot}$ and $10^{13.6}\ {\rm M}_{\odot}$, respectively, we require large fraction of contamination to see the effect of misclassified centrals compared to that for the MW-stellar mass bin. More precisely, we find that the effect of misclassified centrals dominates over the central galaxy emission at $R_{500c}$ with satellite contamination fractions, $f_{\rm cont,\ sat}$, of $0.01,0.14$, and $0.77$ for the stellar mass bins $10^{10.5-11}\ {\rm M}_{\odot}$, $10^{11-11.25}\ {\rm M}_{\odot}$, and $10^{11.25-11.5}\ {\rm M}_{\odot}$, respectively. Hence, the contamination fractions probed here ($f_{\rm cont,\ sat}=0.01,0.1,$ and $0.3$), which are closer to the realistic $f_{\rm cont,\ sat}$ in observations, do not significantly affect the stellar mass bins $10^{11-11.25}$, and $10^{11.25-11.5}$, respectively.
We now discuss how to observationally mitigate the effect of satellites misclassified as centrals using X-ray information. We investigate this issue here under the assumption that the observational galaxy sample is accurately classified into centrals and satellites; in such a case, the effect of misclassified centrals is dominated by instrument resolution, i.e., only satellite objects that are farther from the central than the distance corresponding to the instrument resolution can be easily disregarded as misidentified centrals. However, we must also note that using the X-ray information to mitigate the effect of satellites (based on their proximity to a central) requires the individual halos to be detected in X-rays, which, given the faintness of the CGM emission, is currently not viable for MW-mass halos with survey instruments like eROSITA (see further discussion in Sec. 6.3). For more massive (M31-like) halos, which have higher chances of detection in X-rays, we study how such segregation of misclassified centrals based on their proximity to a central in the sky would impact the mean X-ray surface brightness profile from misclassified centrals. We find that the mean X-ray profile from M31-mass satellite galaxies within $15"$ of a central galaxy drops by $25\%$ at $R_{500c}$ as opposed to the case where we consider all the satellite galaxies as misclassified centrals, as shown in Fig. 11. Therefore, such a technique for mitigating satellite contamination reduces the flattening of the mean X-ray surface brightness profile from satellites. However, an important caveat when selecting centrals without satellites in the vicinity is that such an experiment would give a biased result in the detected CGM itself, where one would be probing the CGM of galaxies only in isolated environments.
@@
Given these findings, we conclude that the contribution of satellites in MW-like galaxy samples containing satellites must not be neglected in stacking analysis. Observationally, to disentangle the contamination of the misclassified centrals, we must fit the profiles contributing to the intrinsic central galaxy jointly with the component capturing the effect of misclassified centrals for a given fraction of satellite contamination.
Figure 11: Effect of further selection on the satellites misclassified as central galaxies on the X-ray surface brightness profile. The total population of satellites misclassified as centrals (solid green line) is divided into satellites within 15” of a central (green dotted line) and satellites farther than 15” from a central (grey dotted line) in projection on the sky. We show that satellites (misclassified as centrals) within 15” of a central galaxy have a steeper profile than the case where we consider all the satellites, thereby reducing the overall flattening of the X-ray surface brightness profile due to misclassified centrals. Contrarily, the satellites farther than 15” from a central have a constant surface brightness profile. The central galaxy profile is shown in purple. The vertical dashed line at $369$ kpc is the mean $R_{\rm 500c}$ of the $680$ galaxies used in the M31 stellar mass bin, and the shaded area corresponds to the minimum and maximum $R_{\rm 500c}$ value of the central galaxies in the M31-mass bin. The shaded uncertainties on the profiles represent the variance obtained by bootstrapping.
6.3 The effect of miscentering: X-ray centre offset from the centre defined by the minimum of the dark matter potential
Here, we study the effect of offset between the observationally defined X-ray centre and the centre defined by the minimum of the halo dark matter potential. The X-ray centre is computed using a luminosity-weighted average of all the counts within the spherical $R_{\rm 200m}$ cubes of the central halo. In this work, we use the input centre or the true centre as the location of the most bound particle within the halo as given by the FoF/subfind algorithm applied on TNG300 to obtain halo catalogues. We use the projected positions on the sky for computing the offset, $\Delta_{\rm X-ray-input}$, and convert these offsets to physical scales using the redshift of the halo. The mean of the $\Delta_{\rm X-ray-input}$ distribution is $70$ kpc, with the $16^{\rm th}-84^{\rm th}$ percentiles ranging between $13-131$ kpc, which is consistent with other works in literature (Seppi et al. 2023; Popesso et al. 2024). The corresponding mean value of the $M_{\rm 200m}$ distribution is at $10^{13.3}{\rm M}_{\odot}$, i.e., within the M31-mass bin, with $16^{\rm th}-84^{\rm th}$ percentiles ranging between $10^{12.4-13.5}{\rm M}_{\odot}$. The mean offset in the three halo mass bins considered in this work, $10^{12.5-13}{\rm M}_{\odot}$, $10^{13-13.5}{\rm M}_{\odot}$, and $10^{13.5-14}{\rm M}_{\odot}$, are $52$ kpc, $66$ kpc, and $79$ kpc, respectively; thereby, showing a correlation of increasing X-ray to true centre offset with increasing halo masses.
A galaxy stacking experiment with an eROSITA-like instrument that uses the dark matter halo potential minima as the centre would be affected by this miscentering effect. Assuming that the X-ray halo is detected, the objects for which the miscentering effect can be corrected are those where the offset between the X-ray and true centre is $\gtrsim 30"$, the resolution of eROSITA. In our lightcone, we report that $\sim 13\%$ of objects have resolvable offsets (i.e., the offset separation $\gtrsim 30"$), which, if classified in $M_{\rm 200m}$ halo mass bins are as follows. For the $10^{12.5-13}{\rm M}_{\odot}$ and $10^{13-13.5}{\rm M}_{\odot}$ mass bins, $13\%$ of objects have resolvable offsets; however, for the more massive $10^{13.5-14}{\rm M}_{\odot}$ mass bin, $17\%$ of objects have resolvable offsets. However, we must note that individual MW mass halos in X-rays with eROSITA are below the detection threshold. Therefore, detecting their X-ray peak is currently unfeasible, but might be possible with future observatories.
@@ 7 Summary
    We present an analytical model (Eq. 7) that well-describes the intrinsic $S_{X}$-profile in LC-TNG300 across the stellar mass bins $10^{10.5-11}{\rm M}_{\odot}$, $10^{11-11.25}{\rm M}_{\odot}$, and $10^{11.25-11.5}{\rm M}_{\odot}$ and halo mass bins of $10^{12.5-13}{\rm M}_{\odot}$, $10^{13-13.5}{\rm M}_{\odot}$, and $10^{13.5-14}{\rm M}_{\odot}$. We provide the best-fitting parameters for the analytical model in Tab. 4 and Tab. 4 for the stellar and halo mass bins explored in this work.
@@
    We present, for the first time, the effect of misclassified centrals in stacking experiments for three stellar mass bins $10^{10.5-11}{\rm M}_{\odot}$, $10^{11-11.25}{\rm M}_{\odot}$, and $10^{11.25-11.5}{\rm M}_{\odot}$. We find that the contaminating effect of the misclassified centrals on the stacked profiles increases when the stellar mass decreases.
@@
    For the MW-like galaxies, we conclude that the contribution of satellites (or misclassified centrals) can not be neglected in stacking analysis (see Fig. 10). In cases where $30\%$, $10\%$, or $1\%$ of the satellites contribute to the total emission of MW-like galaxies, we have shown that they dominate the measured total $S_{X}$ profile at radii $\geq 0.11\times R_{500c}$, $\geq 0.24\times R_{500c},$ and $\geq 1.04\times R_{500c}$, respectively.
  5. 5.
    We report the mean offset between the centre of the halo defined using the X-ray peak, and that obtained from the minimum of the halo potential is $70$ kpc.
Modelling observed CGM profiles and comparing them with simulations is crucial to constrain the different galaxy formation models and to understand how the feedback and physics prescriptions affect the hot CGM profile. Current state-of-the-art cosmological hydrodynamical simulations are calibrated on the stellar mass function and successfully reproduce realistic galaxy populations. Despite this, Davies et al. (2020a) show that EAGLE and IllustrisTNG predict different total gas mass fractions, affecting the observed CGM properties at MW-masses. Similarly, Khrykin et al. (2024) shows how the hot gas is sensitive to the different feedback variants within SIMBA. The model and methodology presented here provide the machinery to compare the measured hot CGM profiles among current simulations in future works.
Future X-ray missions, on the observation side, like Athena (Nandra et al. 2013), AXIS (Mushotzky et al. 2019), HUBS (Cui et al. 2020) will push our current detection limits and provide us with (i) the spatial resolution to reach higher redshifts, (ii) better quantify point-like source contamination within the hot CGM, (iii) the spectral resolution to allow disentangling components via spectral fitting and (iv) the grasp to reach even fainter surface brightness levels. This work, focusing on modelling the projection effects, is a step towards exploiting the information provided by the next-generation telescopes to better understand the intrinsic hot CGM emission. The prospects of this framework would be to extend it to interpret hot X-ray CGM measurements in stacking experiments by accounting for all projection effects, i.e., not only the impact due to the local environment of the halo and the effect due to misclassified centrals but also the emission from other X-ray sources such as AGN and XRB.
@@ Acknowledgements.
SS would like to thank Fulvio Ferlito, Matteo Guardiani, Christian Garrel, Emre Bahar, Jeremy Sanders, and the anonymous referee for all the useful scientific discussions. G.P. acknowledges financial support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program ”Hot Milk” (grant agreement No. 865637) and support from Bando per il Finanziamento della Ricerca Fondamentale 2022 dell’Istituto Nazionale di Astrofisica (INAF): GO Large program and from the Framework per l’Attrazione e il Rafforzamento delle Eccellenze (FARE) per la ricerca in Italia (R20L5S39T9). Computations were performed on the HPC system Raven at the Max Planck Computing and Data Facility. We acknowledge the project support by the Max Planck Computing and Data Facility.